Prove by means of slope that the three points (3,1), (6,0) and (4,4)
are the vertices of a right triangle.

Answer by math_helper(2461)

Label the points:
A @ (3,1)
B @ (6,0)
C @ (4,4)

Slope_AB = (0-1)/(6-3) = -1/3
Slope_AC = (4-1)/(4-3) = 3

We don't need to go any further. Slope_AB = -1/Slope_AB so we know immediately line segment AB is perpendicular to line segment AC and ABC is a right triangle.

(Lines with slopes m1 and m2 are perpendicular if their slopes are related by m2 = -1/m1)